Use the given information to find the number of degrees of freedom, the critical values chi Subscript Upper L Superscript 2χ2L and chi Subscript Upper R Superscript 2χ2R , and the confidence interval estimate of sigmaσ. It is reasonable to assume that a simple random sample has been selected from a population with a normal distribution. Nicotine in menthol cigarettes 98 % confidence; nequals=23 , sequals=0.27 mg.
Answer #1

Let n = sample size = 23

degrees of freedom = n - 1 = 23 - 1 = 22

confidence level = c = 0.98

level of significance = alpha = 1 - c = 1- 0.98 = 0.02

therefore, alpha /2 = 0.02/2 = 0.01

Let's use excel to find the critical chi-square values:

2R = "=CHIINV(0.01,22)" = 40.2894

chi ^{2}_{L} = "=CHIINV(1-0.01,22)" = 9.5425

It is given that s = 0.27

Formula of confidence interval for sigma is as follows:

S k

Plug the values in this formula we get:

-<σ< 0.27* * 0.1995 < σ < 0.4100 40.2894 9.5425

Therefore lower limit = 0.1995

and the upper limit = 0.4100